an object of maas m1 moving with speed V  collide with anorther object of mass m2 at rest stick to it. find the impulse imparted to second object?

Grade:12th pass

2 Answers

Vikas TU
7 years ago
From the momentum Conservation,
m1v + 0 = (m1+m2)v’
v’ = m1v/(m1+m2) m/s.
 
J = change in momentum ( for second object.)
that is = > Pf – Pi
=> (m2)v’ - m2*0 
=> m1*m2*v/(m1+m2) N-m.
Piyush Kumar Behera
7 years ago
Let v1 be the combined velocity of both the bodies after collision.
Momentum of the system before collision=m1v
momentum of the system after collision =(m1+m2)v1
As the total momentum is conserved,so 
m1v=(m1+m2)v1
v1=m1v/(m1+m2)
Momentum of m2 before collision=0
momentum of m2 after collision =m2v1=m1m2v/(m1+m2)
impulse imparted=final momentum – initial momentum
=m1m2v/(m1+m2)-0=m1m2v/(m1+m2)
so the impulse imparted to the second object is m1m2v/(m1+m2)